OpenStudy (secret-ninja):

Which set of line segments with the given lengths form a right triangle? A. 9, 27, 81 B. 10, 24, 26 C. 13, 15, 30 D. 15, 18, 23

OpenStudy (paki):

FOR PART A.... here... a = 9... b = 27... and c = 81... using phytagoras theorem.... a^2 + b^2 = c^2 9^2 + 27^2 = 81^2 81 + 729 = 6561 810 = 6561 ==> as both sides are not equal.... similarly, do this procedure for the remaining... which side you got as balanced, that will be the right angle triangle sides.... @Secret-Ninja

OpenStudy (secret-ninja):

Wait, so I use this equation, and the number I get... wait, what? I dont understand this...

OpenStudy (shubhamsrg):

You actually do not have to do this much calculation. You can always eliminate choices A and C since these 2 fail the basic criteria which is required for a triangle to be possible, i.e. sum of 2 sides should be greater than the third. Just do what @paki told you for options B and D.

OpenStudy (shubhamsrg):

Square each side and check if the largest of the squares is the sum of the squares of other 2. If this is true, then they form a right angled triangle, else not.

OpenStudy (paki):

yup this method is also true....

OpenStudy (secret-ninja):

oh i get it now, thanks!

OpenStudy (secret-ninja):

thanks paki but he made it sound easier XD lol

OpenStudy (shubhamsrg):

Glad to help. :) Paki's method was universal though and can be applied always ^^

OpenStudy (paki):

hmmm always pleasure :-)